An overseas phone call is charged at one rate (a fixed amount) for the first minute and at a different rate for each additional minute. If a 7 minute call costs $10, and a 4 minute call costs $6.40, find each rate.
Answer by mananth(16946)

An overseas phone call is charged at one rate (a fixed amount) for the first minute and at a different rate for each additional minute. If a 7 minute call costs $10, and a 4 minute call costs $6.40, find each rate.

Seven minute call
let the first minute cost be x
balance will be 10-x for the balance 6 minutes
rate /minute = 10-x / 6
SFour minute call
let the first minute cost be x
balance will be 6.4 -x for the balance 3 minutes
rate /minute = 6.4-x / 3
10-x / 6 =6.4-x / 3
30-3x= 38.4 - 6.4x
3x= 8.4
x=2.8
if x= 2.8 then the balance minutes will be charged at $ 1.2 /minute.